### v1

@@ Abstract
While Multimodal Large Language Models (MLLMs) have experienced significant advancement on visual understanding and reasoning, their potentials to serve as powerful, flexible, interpretable, and text-driven models for Image Quality Assessment (IQA) remains largely unexplored. In this paper, we conduct a comprehensive and systematic study of prompting MLLMs for IQA. Specifically, we first investigate nine prompting systems for MLLMs as the combinations of three standardized testing procedures in psychophysics (i.e., the single-stimulus, double-stimulus, and multiple-stimulus methods) and three popular prompting strategies in natural language processing (i.e., the standard, in-context, and chain-of-thought prompting). We then present a difficult sample selection procedure, taking into account sample diversity and uncertainty, to further challenge MLLMs equipped with the respective optimal prompting systems. We assess three open-source and one close-source MLLMs on several visual attributes of image quality (e.g., structural and textural distortions, color differences, and geometric transformations) in both full-reference and no-reference scenarios. Experimental results show that only the close-source GPT-4V provides a reasonable account for human perception of image quality, but is weak at discriminating fine-grained quality variations (e.g., color differences) and at comparing visual quality of multiple images, tasks humans can perform effortlessly.
@@
Image quality assessment Multimodal large language models Model benchmarking
Figure 1: Illustration of several visual attributes of image quality in our experiments.
@@ 1 Introduction
The evolution of Large Language Models (LLMs) has marked a significant milestone in the field of Artificial Intelligence (AI) [53, 49, 6, 2]. The underlying idea of scaling the model size and training data [20] has rendered LLMs the abilities to perform various Natural Language Processing (NLP) tasks with unprecedented levels of accuracy [2].
In the midst of these developments, a particularly promising offshoot has emerged in the form of Multimodal LLMs (MLLMs). These advanced models have taken the capabilities of LLMs a step further by incorporating visual data alongside text [31, 64, 13, 4, 46, 61]. MLLMs typically integrate visual data via Vision Transformers (ViTs) [14] for feature extraction, attention mechanisms [50] for visual-textual relationship modeling, and fusion modules [3, 26, 17] to merge the two modalities. These techniques enable MLLMs to process both text and image data holistically, extending the application scenario of LLMs.
@@
  • Improved robustness. Being sequential in nature, MLLMs do not rely on alignment of images for quality assessment, making them easily robust to perturbations that are imperceptible to the human eye, such as mild geometrical transformations [34] and texture substitution [11].
@@
  • Interpretability. MLLMs generate descriptive text rather than merely providing a numerical score [67]. This text output allows for more detailed, contextually rich, and human-like quality assessment, making MLLMs valuable for IQA model-in-the-loop image enhancement [29].
Previous work [55, 56, 67] focuses primarily on establishing IQA datasets with human quality descriptions to benchmark MLLMs in terms of quality question answering, rating, and reasoning. Appealing at first glance, these studies may suffer from several limitations. First, how to properly instruct human subjects to supply detailed and balanced descriptions of image quality and other relevant visual information is highly nontrivial. Second, image quality is a perceptual quantity with subjective biases. How to screen outlier descriptions and aggregate valid but relatively inconsistent descriptions are overlooked. Third, comparing MLLM outputs to reference descriptions is complex, and remains an intriguing and challenging problem in language modeling [28, 1]. After all, the field of IQA has a rich history, and numerous established human-rated image quality datasets are readily accessible for evaluating this perceptual aspect of MLLMs. Closest to ours, Zhu et al. [73] tested the pairwise quality comparison capability of MLLMs in a constrained context.
In this work, we further pursue this direction, and conduct a comprehensive and systematic study of prompting MLLMs for IQA, as opposed to optimizing individual text prompts. We first explore nine prompting systems for MLLMs, combining standardized testing procedures in psychophysics (i.e., the single-stimulus, double-stimulus, and multiple-stimulus methods) with popular prompting strategies in NLP (i.e., the standard, in-context, and chain-of-thought prompting). Interestingly, we discover that the optimal prompting system varies between open-source [31, 64, 13] and closed-source MLLMs. To further challenge MLLMs, we propose a computational procedure to select difficult samples using top-performing IQA expert models as proxies [60, 71, 11, 69, 57, 8], while taking into account sample diversity and uncertainty. Under both Full-Reference (FR) and No-Reference (NR) settings, we experiment with three open-source and one close-source MLLMs across several visual attributes of image quality, including structural and textural distortions, color differences, and geometric transformations (see Fig. 1). Experimental results show that even the most advanced MLLM (i.e., the proprietary GPT-4V) still has room for improvement in modeling the incredibly sophisticated adaptable HVS for IQA.
@@ 2.1 Expert Models for IQA
IQA can be roughly divided into two categories: FR- and NR-IQA. FR-IQA models are preferred in situations where the unimpaired reference image is available, such as in many low-level vision tasks. Representative design philosophies for FR-IQA range from measuring error visibility (e.g., the mean squared error (MSE)), structural similarity (e.g., the SSIM index [52]), and mutual information (e.g., the VIF measure [43]) to (deep) learning-based methods (e.g., the LPIPS [69] and DISTS metrics [11]) and fusion-based approaches (e.g., VMAF [48]). Almost all FR-IQA models depend on the proper alignment of the reference and test images to execute co-located comparison at the pixel, patch, or feature level. Consequently, these models may struggle to capture the robustness of the HVS to mild geometric transformations and texture resampling. Additionally, these models frequently disregard or give a superficial treatment of color information, leading to a poor account for perceptual color differences [51, 9].
NR-IQA models [37, 70, 60, 21, 71, 59, 63] are the most challenging yet widely applicable as they are designed to evaluate image quality without any reference. Regardless of whether they are knowledge-driven [37] or data-driven [70, 71], NR-IQA models rely heavily on the training data, which often demonstrate weak generalization when faced with a broader type of images and distortions.
Apart from accuracy and generalization, expert IQA models also fall short in terms of flexibility and interpretability. This is primarily because they reduce the evaluation of image quality to a numerical score, failing to leverage other formats of inputs, and output other detailed and perceptually relevant information.
@@ 2.2 MLLMs for IQA
Previous work has been centered on benchmarking and refining MLLMs for IQA. The pioneering datasets, DepictQA [67] and Q-Bench [55], have initiated the assessment of MLLMs in quality rating and reasoning under the FR and NR conditions, respectively. DepictQA employs a double-stimulus method, prompting MLLMs to compare a pair of images against a standard reference. In contrast, Q-Bench takes a single-stimulus approach, assigning a numerical score based on the classification of image quality as “poor” or “good”. Notably, these investigations reveal that, except for the proprietary GPT-4V [61], open-source MLLMs like LLaVA [31], Kosmos-2 [41], and MiniGPT-4 [72] show limited success in replicating human perception of image quality. Efforts to enhance open-source MLLMs through instruction tuning have been made [56], focusing on the prediction of subsequent tokens. Yet, such models often revert to generating template-like quality descriptions, lacking the level of flexibility we are looking for. In this work, we take a step back and conduct a comprehensive and systematic exploration of prompting techniques for MLLMs on existing human-rated IQA datasets.
@@
3 Systematic Prompting MLLMs for IQA
In this section, we introduce several prompting systems for MLLMs in the context of IQA, followed by describing the difficult sample selection procedure.
3.1 Systematic Prompting from Psychophysics
Although MLLMs and the HVS (or more generally the human brain) operate in different domains, they have fascinating similarities, especially in representational hierarchy [5], visual-textual integration, contextual reasoning [22], and adaptation [12]. Thus, it is natural to employ standardized psychophysical testing procedures as the systematic prompting methods for IQA.
Single-stimulus Method. As depicted in Figure 2 (a), given a test image $x\in\mathcal{X}$, the FR single-stimulus method [^1] This is also known as the double-stimulus impairment rating by treating the reference image as a second visual stimulus. suggests MLLMs to output a quality score, $q(x)\in[0,100]$, based on its perceptual difference to the corresponding reference image $y\in\mathcal{Y}$. Here, a larger number indicates higher quality. In contrast, the NR single-stimulus method derives the quality score solely from the perceptually important visual attributes of the test image $x$. Generally, the single-stimulus method is scalable and straightforward to implement for MLLMs, but may not capture relative quality accurately and may exhibit context dependency (e.g., the presentation order).
@@ 3.1 Systematic Prompting from Psychophysics
Double-stimulus Method. Also known as two-alternative forced choice (2AFC) and paired comparison, the double-stimulus method first samples a pair of images either uniformly or actively [62] from a set of images $\mathcal{X}=\{x_{i}\}^{M}_{i=1}$, where $M$ is the number of images. In our implementation, MLLMs are not forced to select between the two alternatives with higher quality. Instead, they have a third option: the quality of the two images is comparable. Figure 2 (b) illustrates the FR and NR double-stimulus methods. In the FR scenario, the pair of images are constrained to share the same underlying visual content but differ in types and intensities of distortions. Additionally, a reference image is supplied to facilitate quality comparison. In the NR scenario, the pair of images can be of different content, without supplying any reference image. Upon constructing the pairwise preference matrix $C\in\mathbb{R}^{M\times M}$, where
$C_{ij}=\begin{cases}\#\text{ times }x_{i}\text{ preferred over }x_{j}&\text{if}i\neq j\\ 0&\text{if }i=j,\end{cases}$ (1)
we adopt the maximum a posteriori estimation to aggregate pairwise rankings, under the Thurstone’s Case V assumption [47]. This amounts to solving the following convex optimization problem:
[…]
subject to $\sum_{i}\mu_{i}=0,$
where $\mu=[\mu_{1},\mu_{2},\dots,\mu_{M}]$ are the global quality scores for the $M$ images and $\mathrm{\Phi}(\cdot)$ is the standard normal cumulative distribution function. Alternative pairwise ranking aggregation methods can be employed, such as the Thurstone-Mosteller least squares, the Perron-Frobenius ranking, and the ELO system [15]. We find empirically that these approaches all yield reliable and consistent ranking outcomes. Arguably the double-stimulus method is considered more reliable than the single-stimulus counterpart when gathering human opinions of image quality, despite its $O(M^{2})$ sample complexity. It remains to be seen whether such reliability extends to prompting MLLMs.
@@
Multiple-stimulus Method. As shown in Figure 2 (c), the multiple-stimulus method proposes a more efficient approach to gather partial rankings by simultaneously presenting a set of $L$ images to MLLMs for ranking. Just like the double-stimulus method, in the FR scenario, we limit the list of $L$ images to have identical visual content and supply the corresponding reference image. This restriction and the presence of the reference image are not assumed in the NR scenario. The listwise ranking of $L$ images yields $\binom{L}{2}$ pairwise rankings, which are used to form the pairwise preference matrix $C$ in Eq. (1), followed by ranking aggregation using Eq. (2).
Figure 3 gives the detailed text prompts to implement the single-stimulus, double-stimulus and multiple-stimulus methods for GPT-4V in the NR-IQA scenario. It is crucial to acknowledge that our priority is not to optimize individual textual prompts at the token level, but to explore the impact systematic prompting methods on IQA.
@@
3.2 Systematic Prompting from NLP
MLLMs evolve from LLMs, thereby inherently accommodating systematic prompting strategies in NLP.
Chain-of-thought Prompting. As a simple yet powerful approach to encouraging step-by-step reasoning, chain-of-thought prompting [54] has proven to be highly beneficial for a range of NLP tasks. In the same vein, we request MLLMs to detail the perceptual quality of the test image(s) by examining various visual attributes (and comparing to the reference image(s) when available), before integrating the analysis results into overall quality estimate(s) (see Figure 3 (b)).
@@ 3.2 Systematic Prompting from NLP
By integrating prompting strategies derived from psychophysics and NLP, we end up with nine prospective prompting systems for testing. Given the variations in model architectures, training data, optimization pipelines, and alignment paradigms [38, 74, 39], it is anticipated that different MLLMs may function optimally with different prompting systems.
@@ 3.3 Computational Procedure for Difficult Sample Selection
Inference with MLLMs tends to be slow and costly, making it impractical to evaluate MLLMs on the full IQA datasets paired with the proposed nine prompting systems. We describe a computational procedure to pinpoint a smaller set of the most informative samples that meet three key criteria.
First, they should be difficult, with a high likelihood of causing MLLMs to err. The Group MAximum Differentiation (gMAD) competition [36] has been used before for difficult sample selection, but it is not applicable here due to the violation of affordable inference assumption. We thus utilize efficient expert IQA models as proxies, and identify their failures by maximizing the MSE between model predictions and human quality scores, as a form of “black-box” attacks [40] on MLLMs. Second, the selected samples should be diverse to highlight various aspects of MLLMs’ potential weaknesses in IQA. Third, they should be consistent with small variations in individual quality scores. That is, samples that have inconsistent or uncertain human ratings are less meaningful to assess MLLMs. Putting together, we have the following optimization problem in the FR scenario:
@@
where $\mathcal{Y}^{\star}_{n-1}=\{y^{\star}_{n^{\prime}}\}_{n^{\prime}=1}^{n-1}$ contains the selected reference images in the previous $n-1$ iterations from $\mathcal{Y}$, the set of all reference images. $\mathcal{X}_{y}$ includes all distorted images that originate from $y$. $d_{w}(\cdot,\cdot)$ denotes the expert FR-IQA model, parameterized by $w$. $q(\cdot)$ and $\sigma(\cdot)$ are respectively the human quality score and standard deviation. $\mathrm{Div}(\cdot)$ is the point-to-set measure, quantifying the added diversity of $y$ to $\mathcal{Y}^{\star}_{n-1}$. $\lambda$ is a parameter, trading off the variance-normalized squared error and the diversity measure. $\epsilon$ is a small positive constant to avoid any potential division by zero.
@@
where $q_{w}(\cdot)$ denotes the expert NR-IQA model. Figure 4 shows the comparison between difficult sample selection with and without variance normalization (as a measure of sample uncertainty) under the same level of sample diversity. We find that variance-normalized sampling induces a noticeable shift towards zero in the empirical distribution of sample standard deviations, suggesting an enhanced level of human consistency in assessing image quality.

### v2

@@ Abstract
While Multimodal Large Language Models (MLLMs) have experienced significant advancement in visual understanding and reasoning, their potential to serve as powerful, flexible, interpretable, and text-driven models for Image Quality Assessment (IQA) remains largely unexplored. In this paper, we conduct a comprehensive and systematic study of prompting MLLMs for IQA. We first investigate nine prompting systems for MLLMs as the combinations of three standardized testing procedures in psychophysics (i.e., the single-stimulus, double-stimulus, and multiple-stimulus methods) and three popular prompting strategies in natural language processing (i.e., the standard, in-context, and chain-of-thought prompting). We then present a difficult sample selection procedure, taking into account sample diversity and uncertainty, to further challenge MLLMs equipped with the respective optimal prompting systems. We assess three open-source and one closed-source MLLMs on several visual attributes of image quality (e.g., structural and textural distortions, geometric transformations, and color differences) in both full-reference and no-reference scenarios. Experimental results show that only the closed-source GPT-4V provides a reasonable account for human perception of image quality, but is weak at discriminating fine-grained quality variations (e.g., color differences) and at comparing visual quality of multiple images, tasks humans can perform effortlessly.
@@
Image quality assessment Multimodal large language models Model comparison
Figure 1: Illustration of visual attributes of image quality in our experiments.
@@ 1 Introduction
The evolution of Large Language Models (LLMs) has marked a significant milestone in the field of Artificial Intelligence (AI) [51, 47, 6, 2]. The underlying idea of scaling the model size and training data [19] has rendered LLMs the abilities to perform various Natural Language Processing (NLP) tasks with unprecedented levels of accuracy.
In the midst of these developments, a particularly promising offshoot has emerged in the form of Multimodal LLMs (MLLMs). These advanced models have taken the capabilities of LLMs a step further by incorporating visual data alongside text [29, 62, 13, 4, 43, 59]. MLLMs typically integrate visual data via Vision Transformers (ViTs) [14] for feature extraction, attention mechanisms [48] for visual-textual relationship modeling, and connector modules [3, 24, 16, 45] to merge the two modalities. These techniques enable MLLMs to process both text and image data holistically, extending the application scenario of LLMs.
@@
  • Improved robustness. Being sequential in nature, MLLMs do not rely on the alignment of images for quality assessment, making them easily robust to perturbations that are imperceptible to the human eye, such as mild geometrical transformations [31] and texture resampling [11].
@@
  • Interpretability. MLLMs generate descriptive text rather than merely providing a numerical score [66]. This text output allows for more detailed, contextually rich, and human-like quality assessment, making MLLMs valuable for IQA model-in-the-loop image processing [27].
Previous work [53, 54, 66] focuses primarily on establishing IQA datasets with human quality descriptions to benchmark MLLMs in terms of quality question answering, rating, and reasoning. Appealing at first glance, these studies may suffer from several limitations. First, how to properly instruct human subjects to supply detailed and balanced descriptions of image quality and other relevant visual information is highly nontrivial. Second, image quality is a perceptual quantity with subjective biases. How to screen outlier descriptions and aggregate valid but relatively inconsistent descriptions are overlooked. Third, comparing MLLM outputs to reference descriptions is complex, and remains an intriguing and challenging problem in language modeling [26, 1]. After all, the field of IQA has a rich history, and numerous established human-rated image quality datasets are readily accessible for evaluating this perceptual aspect of MLLMs.
In this work, we conduct a comprehensive and systematic study of prompting MLLMs for IQA. We first explore nine prompting systems for MLLMs, combining standardized testing procedures in psychophysics (i.e., the single-stimulus, double-stimulus, and multiple-stimulus methods) with popular prompting strategies in NLP (i.e., the standard, in-context, and chain-of-thought prompting). To further challenge MLLMs, we propose a computational procedure to select difficult samples using top-performing IQA expert models as proxies [58, 70, 11, 68, 55, 8], while taking into account sample diversity and uncertainty. Under both Full-Reference (FR) and No-Reference (NR) settings, we experiment with three open-source and one closed-source MLLMs across several visual attributes of image quality, including structural and textural distortions, geometric transformations, and color differences (see Fig. 1).
Our experimental results highlight three key takeaways. First, different MLLMs require different prompting systems to perform optimally. This implies the need for a re-examination of the recent progress in MLLMs for IQA, particularly in comparison to GPT-4V [56, 66, 65]. Second, aided by the proposed difficult sample selection method, we demonstrate that there is still ample room for improving MLLMs (including GPT-4V) for IQA, especially in fine-grained quality discrimination and multiple-image quality analysis. Third, we argue that directly fine-tuning open-source MLLMs on datasets with image quality descriptions may not be effective due to the risk of catastrophic forgetting of models’ general abilities in visual reasoning and understanding. The encouraging results from the chain-of-thought prompting suggest that IQA should be integrated into a broader and higher-level task, making use of additional physical, geometrical, and semantic information to infer image quality.
@@ 2.1 Expert Models for IQA
IQA can be divided into two categories: FR- and NR-IQA. FR-IQA models are preferred in situations where the undistorted reference image is available. Representative design philosophies for FR-IQA range from measuring error visibility (e.g., the mean squared error (MSE)), structural similarity (e.g., the SSIM index [50]), and mutual information (e.g., the VIF measure [40]) to (deep) learning-based methods (e.g., the LPIPS [68] and DISTS metrics [11]) and to fusion-based approaches (e.g., VMAF [46]). Almost all FR-IQA models depend on the proper alignment of the reference and test images to execute co-located comparison at the pixel, patch, or feature level. Consequently, these models may struggle to capture the robustness of the HVS to mild geometric transformations and texture resampling. Additionally, these models often give a superficial treatment of color information, leading to a poor account for perceptual color differences [49, 9].
NR-IQA models [34, 69, 58, 20, 70, 57, 61] are more challenging yet widely applicable as they evaluate image quality without any reference. Whether they are knowledge-driven [34] or data-driven [69, 70], NR-IQA models rely heavily on human-annotated training data. This reliance leads to weak generalization when applied to a wider range of images and distortions.
Apart from accuracy and generalization, expert IQA models also fall short in terms of flexibility and interpretability. This is primarily because they summarize image quality using a numerical score, missing the opportunity to leverage other input formats and to provide more detailed and perceptually relevant outputs.
@@ 2.2 MLLMs for IQA
Previous work has been centered on benchmarking and fine-tuning MLLMs for IQA. The pioneering datasets, DepictQA [66] and Q-Bench [53], have initiated the assessment of MLLMs in quality rating and reasoning under the FR and NR conditions, respectively. DepictQA employs a double-stimulus method, prompting MLLMs to compare a pair of images against the reference. Q-Bench takes a single-stimulus approach, computing a numerical score based on the classification of image quality as “poor” or “good”. These investigations reveal that, except for the proprietary GPT-4V [59], open-source MLLMs like LLaVA [29], Kosmos-2 [38], and MiniGPT-4 [71] show limited success in replicating human perception of image quality. Efforts to enhance open-source MLLMs through instruction tuning have been made [54]. Yet, such models often revert to generating template-like quality descriptions, lacking the level of flexibility we are looking for. In this work, we take a step back and conduct a comprehensive and systematic exploration of prompting techniques for MLLMs on existing human-rated image quality datasets.
3 Prompting MLLMs for IQA
In this section, we introduce several IQA prompting systems for MLLMs, followed by the description of the difficult sample selection procedure.
3.1 Prompting Strategies from Psychophysics
Although MLLMs and the HVS (or more generally the human brain) operate in different domains, they have fascinating similarities, especially in representational hierarchy [5], visual-textual integration, contextual reasoning [21], and adaptation [12]. Thus, it is natural to employ standardized psychophysical testing procedures as prompting strategies for IQA.
Single-stimulus Method. As depicted in Figure 2 (a), given a test image $x\in\mathcal{X}$, the FR single-stimulus method [^1] This is also known as the double-stimulus impairment rating by treating the reference image as a second visual stimulus. suggests MLLMs to output a quality score, $q(x)\in[0,100]$, based on its perceptual distance to the corresponding reference image $y\in\mathcal{Y}$. Here, a larger $q$ indicates higher quality. The NR single-stimulus method derives the quality score solely from the test image $x$. Generally, the single-stimulus method is scalable and straightforward to implement for MLLMs, but may not capture relative quality accurately.
@@ 3.1 Prompting Strategies from Psychophysics
Double-stimulus Method. Also known as two-alternative forced choice (2AFC) and paired comparison, the double-stimulus method first samples a pair of images either uniformly or actively [60] from a set of images $\mathcal{X}=\{x_{i}\}^{M}_{i=1}$, where $M$ is the total number of images. In our implementation, MLLMs are not forced to select between the two alternatives with higher quality. Instead, they have a third option to indicate that the quality of the two images is comparable. Figure 2 (b) illustrates the FR and NR double-stimulus methods. In the FR scenario, the pair of images are constrained to share the same underlying visual content but differ in distortion types and intensities. The reference image is accompanied to facilitate quality comparison. In the NR scenario, the pair of images can be of different content, without supplying any reference image. Upon constructing the pairwise preference matrix $C\in\mathbb{R}^{M\times M}$, where $C_{ij}$ records the number of times $x_{i}$ is preferred over $x_{j}$ for $i\neq j$, we adopt the maximum a posteriori estimation to aggregate pairwise rankings under the Thurstone’s Case V model [44]. Arguably, the double-stimulus method is considered more reliable than the single-stimulus one when gathering human opinions of image quality, despite its $O(M^{2})$ sample complexity. It remains to be seen whether such reliability extends to prompting MLLMs.
@@
Multiple-stimulus Method. As shown in Figure 2 (c), the multiple-stimulus method presents a more efficient approach to gathering partial rankings by simultaneously presenting a set of $L$ images to MLLMs. Just like the double-stimulus method, in the FR scenario, we limit the list of $L$ images to have identical visual content and supply the corresponding reference image. This restriction and the presence of the reference image are not assumed in the NR scenario. The listwise ranking of $L$ images yields $\binom{L}{2}$ pairwise rankings, which are used to form the pairwise preference matrix $C$, followed by ranking aggregation. Figure 3 gives the detailed text prompts to implement the single-stimulus, double-stimulus, and multiple-stimulus methods for GPT-4V in the NR-IQA scenario.
3.2 Prompting Strategies from NLP
@@
Chain-of-thought Prompting. As a simple yet powerful approach to encouraging step-by-step reasoning, chain-of-thought prompting [52] has proven beneficial for a range of NLP tasks. In the same vein, we request MLLMs to detail the perceptual quality of the test image(s) by examining various visual attributes (and comparing them to the reference image(s) when available), before integrating the analysis results into the overall quality estimate(s).
@@
By integrating prompting strategies derived from psychophysics and NLP, we end up with nine candidate prompting systems. Given the variations in training data, model architectures, optimization pipelines, and alignment strategies [35, 73, 36], it is anticipated that different MLLMs may function optimally with different prompting systems.
@@ 3.3 Computational Procedure for Difficult Sample Selection
Inference with MLLMs tends to be slow and costly, making it impractical to evaluate MLLMs on the full IQA datasets, each paired with the nine candidate prompting systems. We describe a computational procedure to pinpoint a smaller set of the most informative testing samples with three desired properties.
First, they should be difficult, with a high likelihood of causing MLLMs to err. The Group MAximum Differentiation (gMAD) competition [33] has been used before for difficult sample selection, but it is not applicable here due to the violation of the affordable inference assumption. We thus utilize efficient expert IQA models as proxies, and identify their failures by maximizing the MSE between model predictions and human quality scores as a form of “black-box” attacks [37] on MLLMs. Second, the selected samples should be diverse to highlight various aspects of MLLMs’ potential weaknesses in IQA. Third, they should be consistent with the majority of human judgments with small variations. Putting together, we formulate the following optimization problem in the FR scenario:
@@
where $\mathcal{Y}^{\star}_{n-1}=\{y^{\star}_{n^{\prime}}\}_{n^{\prime}=1}^{n-1}$ contains the selected reference images in the previous $n-1$ iterations from $\mathcal{Y}$, the set of all reference images. $\mathcal{X}_{y}$ includes all distorted images that originate from $y$. $d_{w}(\cdot,\cdot)$ denotes a generic expert FR-IQA model, parameterized by $w$. $q(\cdot)$ and $\sigma(\cdot)$ are respectively the human quality score and standard deviation. $\mathrm{Div}(\cdot)$ is a point-to-set distance measure, quantifying the added diversity of $y$ to $\mathcal{Y}^{\star}_{n-1}$. $\lambda$ is a parameter, trading off the variance-normalized squared error and the diversity measure. $\epsilon$ is a small positive constant to avoid any potential division by zero.
@@
where $q_{w}(\cdot)$ denotes a generic expert NR-IQA model. Figure 4 shows the comparison between difficult sample selection with and without variance normalization (as a measure of sample uncertainty) under the same level of sample diversity. We find that variance-normalized sampling induces a noticeable shift towards zero in the empirical distribution of sample standard deviations, suggesting an enhanced level of human consistency in assessing image quality.